# Holodeck 2.0: Vision‑Language‑Guided 3D World Generation with Editing

### v1

@@ Abstract
3D scene generation plays a crucial role in gaming, artistic creation, virtual reality and many other domains. However, current 3D scene design still relies heavily on extensive manual effort from creators, and existing automated methods struggle to generate open-domain scenes or support flexible editing. As a result, generating 3D worlds directly from text has garnered increasing attention. In this paper, we introduce Holodeck 2.0, an advanced vision-language-guided framework for 3D world generation with support for interactive scene editing based on human feedback. Holodeck 2.0 can generate diverse and stylistically rich 3D scenes (e.g., realistic, cartoon, anime, and cyberpunk styles) that exhibit high semantic fidelity to fine-grained input descriptions, suitable for both indoor and open-domain environments. Holodeck 2.0 leverages vision-language models (VLMs) to identify and parse the objects required in a scene and generates corresponding high-quality assets via state-of-the-art 3D generative models. It then iteratively applies spatial constraints derived from the VLMs to achieve semantically coherent and physically plausible layouts. Human evaluations and CLIP-based assessments demonstrate that Holodeck 2.0 effectively generates high-quality scenes closely aligned with detailed textual descriptions, consistently outperforming baselines across indoor and open-domain scenarios. Additionally, we provide editing capabilities that flexibly adapt to human feedback, supporting layout refinement and style-consistent object edits. Finally, we present a practical application of Holodeck 2.0 in procedural game modeling, generating visually rich and immersive environments, potentially boosting efficiency.
@@
Figure 1: Examples of stylistically varied 3D scenes generated by Holodeck 2.0. Captions are shortened versions of the original inputs (50–300 words). Holodeck 2.0 faithfully captures fine-grained details from textual descriptions.
Introduction
@@
The ability to generate complex 3D environments from natural language descriptions represents a fundamental challenge at the intersection of computer vision, natural language processing, and computer graphics. As virtual worlds become increasingly central to gaming, film production, architectural visualization, and training embodied AI agents, the demand for efficient, high-quality 3D content creation has never been greater. However, the current practice of 3D scene design still relies predominantly on extensive manual effort by creators. Therefore, the task of generating 3D worlds directly from text or images has attracted widespread attention.
@@
Some current approaches focus on directly generating 3D scenes represented by Neural Radiance Fields (NeRFs) or Gaussian splats from text or images (Poole et al. 2022; Lin et al. 2023; Chung et al. 2023). However, such methods generate scenes that lack separable and manipulable objects, making them difficult to edit directly and limiting their applicability in game design, embodied AI training, etc. Another class of methods focuses on selecting suitable 3D objects from asset datasets and then assembling them into a 3D scene (Paschalidou et al. 2021; Yang et al. 2024; Hu et al. 2024; Aguina-Kang et al. 2024). For example, Holodeck (Yang et al. 2024) uses Large Language Models (LLMs) to parse from the input text the objects required for a scene, then retrieves 3D assets from datasets such as Objaverse (Deitke et al. 2023b) based on Contrastive Language–Image Pre-training (CLIP) (Radford et al. 2021) scores, and finally arranges them into a layout.
@@
Specifically, existing approaches face three critical limitations: (1) asset-based methods are constrained by the quality and diversity of available 3D model libraries, with even large datasets like Objaverse-XL (Deitke et al. 2023a) suffering from inconsistent quality and missing textures; (2) current systems struggle with fine-grained style control, unable to maintain consistent artistic styles (e.g., anime, cyberpunk) across all objects in a scene; and (3) most methods focus exclusively on indoor environments, lacking the flexibility to generate open-world scenarios like fantasy landscapes or futuristic cities.
@@
Recent breakthroughs in vision-language models (VLMs) and 3D generative models present an unprecedented opportunity to address these limitations. VLMs now possess sophisticated spatial reasoning capabilities that can parse complex scene descriptions and generate physically plausible object arrangements. Simultaneously, state-of-the-art 3D generation models like Hunyuan3D can create high-quality, stylistically consistent assets from text or image prompts (Hunyuan3D et al. 2025; Lai et al. 2025). However, these advances have not yet been effectively combined into a unified scene generation framework.
@@
Facilitated by 3D asset generation models and the inference capabilities of VLMs, Holodeck 2.0 can produce high-quality 3D scenes with internal stylistic coherence for both indoor settings (e.g., music studios, playrooms) and open domains (e.g., airship docks, courtyards), as shown in Figure 1. Meanwhile, Holodeck 2.0 can generate the same type of 3D scenes in different styles, such as realistic, cartoon, and steampunk. The objects in Holodeck 2.0 scenes exhibit high quality and closely match their descriptions, for example, “A cartoon-style snail-shaped snack booth offers glowing drinks.” In terms of spatial reasoning, Holodeck 2.0 fully leverages the spatial understanding capabilities of VLMs, iteratively generating and applying spatial constraints to produce layouts that match both textual and visual descriptions.
@@
Beyond research contributions, we demonstrate Holodeck 2.0’s practical applicability through direct integration with commercial game engines. We showcase how a complex museum scene complete with artifacts like the Temple of Dendur, Venus de Milo, and The Thinker—can be generated from detailed text descriptions and seamlessly imported into Unreal Engine with full interactivity. This demonstrates how our system can accelerate 3D content creation workflows for game development and other applications. We demonstrate Holodeck 2.0’s effectiveness through large-scale human evaluations and CLIP-based automated scoring. Across 60 test cases spanning both indoor and open-domain scenes, Holodeck 2.0 consistently outperforms the baseline Holodeck in both human evaluations and CLIP-based automated scoring, with its advantage being particularly pronounced in open-domain scenes.
This paper makes the following contributions:
@@
    We present a unified framework that uniquely combines VLM-guided scene decomposition with generative 3D asset creation, using VLMs across multiple stages—from visual reference generation to object parsing, spatial constraint generation, and iterative layout refinement.
@@
    We demonstrate state-of-the-art performance on both indoor and open-domain scene generation with consistent artistic styles (realistic, cartoon, anime, cyberpunk) through extensive human and automated evaluations.
@@
    We validate practical applicability through integration with commercial game engines and demonstrate significant quality improvements: 110% for indoor scenes and 139% for open-domain scenes compared to prior work.
@@
Figure 2: Overview of Holodeck 2.0. Given a text input, Holodeck 2.0 generates 3D scenes via three modules, with an optional editing module for interactive updates based on human language feedback.
Related Work
Language-Guided 3D Scene Generation
@@
Early work in this area is primarily symbolic and rule‑based, requiring extensive engineering effort and exhibiting very limited generalization ability, for example, Wordseye (Coyne and Sproat 2001) and other systems (Chang, Savva, and Manning 2014; Ma et al. 2018). With the emergence of LLMs and VLMs, open‑vocabulary 3D scene generation has become possible, greatly enhancing the flexibility of scene creation. Holodeck (Yang et al. 2024) employs LLMs to get the object descriptions that should appear in a scene based on the input text, and to generate spatial constraints among those objects. SceneCraft (Hu et al. 2024) also leverages LLMs and VLMs in a similar way, but it generates Blender code and iteratively optimizes it to produce 3D scenes. However, both Holodeck and SceneCraft rely on CLIP (Radford et al. 2021) to retrieve 3D objects from asset libraries such as Objaverse (Deitke et al. 2023b) and TurboSquid (TurboSquid 2024), and the variety and quality of these assets somewhat limit their capability to generate open‑world 3D scenes. Currently, larger or better‑curated 3D asset libraries have emerged (Deitke et al. 2023a; Lin et al. 2025). However, they still suffer from overall low asset quality, missing textures, and limited variety and styles of objects. For 3D scene spatial layout, LayoutGPT (Feng et al. 2023) is a method that directly generates 3D object layout relationships via LLM prompting, but often neglects physical plausibility. LayoutVLM (Sun et al. 2025) is another method that uses VLMs to generate 3D layouts by formulating objective functions that can be differentiably optimized. For 3D scene editing, recent works such as BlenderGPT (BlenderGPT 2023) and BlenderAlchemy (Huang, Yang, and Guibas 2024) have applied LLMs or VLMs to editing, but these models primarily focus on material replacement and perform poorly at placement or layout.
Text-to-3D Diffusion Models
@@
The emergence of diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020; Peebles and Xie 2023; Rombach et al. 2022) marks a significant advancement in the field of generative modeling. Current methods leverage diffusion models to directly generate 3D objects or scenes represented by Neural Radiance Fields (NeRFs) or Gaussian splats from text or images, such as DreamFusion (Poole et al. 2022) and Magic3D (Lin et al. 2023). However, such methods generate scenes that lack discrete, manipulable objects, making them difficult to edit directly and unsuitable for applications like game design or embodied AI training. When focusing solely on the generation of individual 3D assets, diffusion models offer significant advantages. Recent methods have achieved breakthroughs in generating high‑quality 3D assets (Xiang et al. 2025; Zhang et al. 2023, 2024; Zhao et al. 2023; Li et al. 2024), with Hunyuan3D (Hunyuan3D et al. 2025; Lai et al. 2025) being a typical example. It can generate high‑quality, mesh‑based individual 3D assets from text or images.
@@
Holodeck 2.0
Holodeck 2.0 is a vision-language-guided system, facilitated by VLMs and 3D generative models, that can produce and edit diverse, customizable, and stylistically varied open-domain 3D scenes in Blender.
As shown in Figure 2, Holodeck 2.0 systematically generates open-domain 3D scenes from text or images utilizing the following specialized modules: (1) Scene Analysis Module parses the object properties and background information that should appear in the scene from input text or images; (2) Object Generation Module leverages 3D asset generation models to produce 3D assets from quality‑controlled object images; (3) Scene Generation Module iteratively generates semantically and physically coherent object layouts based on spatial constraints; (4) Scene Editing Module automatically applies personalized adjustments to the scene based on human feedback.
In the following, we describe how each module of Holodeck 2.0 operates, specifically illustrating how it coordinates VLMs and 3D asset generation models to accomplish scene generation and editing. We also present its direct application to game design. Detailed information about Holodeck 2.0 is also available in the supplement.
@@
Scene Analysis
The Scene Analysis Module, shown in the first panel of Figure 2, is designed to parse object properties that should appear in the scene from text or images.
Text to Scene Image. Given a text input, the Scene Analysis Module uses VLMs to generate a corresponding reference image. This image generation process is vital for the subsequent creation of uniform‑style assets. Moreover, visual inputs can substantially enhance the spatial awareness of large reasoning models. In Figure 3, we present several examples of scene reference images generated from text.
Scene Image to Object Images. After obtaining the scene image, the Scene Analysis Module employs GPT-o3 (OpenAI 2025) to infer object properties (name, position, rotation, size, visual description, etc.) from the text input and its corresponding scene reference image. Then we prompt the VLMs to generate a clear, transparent frontal image of each object appearing in the scene image. This avoids the issues common to current image segmentation algorithms (Kirillov et al. 2023; Ren et al. 2024), such as inaccurate segmentation, blurry edges, and difficulty handling occluded objects. After generating individual object images, we use GPT-o3 to perform quality control, removing smaller-object images redundantly embedded in larger-object images (e.g., removing the separate faucet when it is already present in the bathtub image). In practice, this approach generates accurate, high‑quality object images from the reference image as shown in Figure 3. Subsequently, we generate the scene’s background image based on the reference image, focusing primarily on the ground texture.
Object Generation
The Object Generation Module, shown in the second panel of Figure 2, utilizes 3D asset generation models to generate 3D assets from quality-controlled object images. For 3D assets generation, image-to-3D is faster and more stable than direct text-to-3D. After obtaining the object images from the Scene Analysis Module, we use Hunyuan3D 2.1 to generate the corresponding 3D asset for each object. It is one of the most powerful open‑source 3D generation models that supports local deployment. In practice, compared to searching for 3D objects in Objaverse, generative models can produce higher‑quality 3D assets that more faithfully match the descriptions (Figure 3), for example, “a plush, L-shaped sectional sofa adorned with colorful throw pillows” and “a cartoon-style donut-shaped bench”.

### v2

@@ Abstract
3D scene generation plays a crucial role in gaming, artistic creation, virtual reality, and many other domains. However, current 3D scene design still relies heavily on extensive manual effort from creators, and existing automated methods struggle to generate open-domain scenes or support flexible editing. To address those challenges, we introduce Holodeck 2.0, an advanced vision-language-guided framework for 3D world generation with support for interactive scene editing based on human feedback. Holodeck 2.0 can generate diverse and stylistically rich 3D scenes (e.g., realistic, cartoon, anime, and cyberpunk styles) that exhibit high semantic fidelity to fine-grained input descriptions, suitable for both indoor and open-domain environments. Holodeck 2.0 leverages vision-language models (VLMs) to identify and parse the objects required in a scene and generates corresponding high-quality assets via state-of-the-art 3D generative models. Then, Holodeck 2.0 iteratively applies spatial constraints derived from the VLMs to achieve semantically coherent and physically plausible layouts. Both human and model evaluations demonstrate that Holodeck 2.0 effectively generates high-quality scenes closely aligned with detailed textual descriptions, consistently outperforming baselines across indoor and open-domain scenarios. Additionally, Holodeck 2.0 provides editing capabilities that flexibly adapt to human feedback, supporting layout refinement and style-consistent object edits. Finally, we present a practical application of Holodeck 2.0 in procedural game modeling to generate visually rich and immersive environments that can boost efficiency in game design.
@@
Figure 1: Examples of stylistically varied 3D scenes generated by Holodeck 2.0. Captions are shortened versions of the original long inputs (50–300 words). Holodeck 2.0 can faithfully capture fine-grained details from textual descriptions.
footnotetext: Equal contribution. Corresponding author.
1 Introduction
@@
Generating complex 3D environments from natural language descriptions is a fundamental challenge at the intersection of computer vision, natural language processing, and graphics. As virtual worlds become increasingly central to gaming, film production, architectural visualization, and training embodied AI agents, the demand for efficient, high-quality 3D content creation has never been greater. However, the current practice of 3D scene design still relies predominantly on extensive manual effort by creators. Therefore, the task of generating 3D worlds directly from text or images has attracted widespread attention.
@@
Some current approaches focus on directly generating 3D scenes represented by Neural Radiance Fields (NeRFs) or Gaussian splats from text or images [poole2022dreamfusion, lin2023magic3d, chung2023luciddreamer]. However, such methods generate scenes that lack separable and manipulable objects, making them difficult to edit directly and limiting their applicability in game design, embodied AI training, etc. Another class of methods focuses on selecting suitable 3D objects from asset datasets and then assembling them into a 3D scene [paschalidou2021atiss, yang2024Holodeck, hu2024scenecraft, aguina2024openuniverse]. For example, Holodeck [yang2024Holodeck] uses Large Language Models (LLMs) to parse from the input text the objects required for a scene, then retrieves 3D assets from datasets such as Objaverse [deitke2023objaverse] based on visual similarities using CLIP [radford2021clip] scores, and finally arranges them into a layout.
@@
Specifically, existing approaches face three critical limitations: (1) asset-based methods are constrained by the quality and diversity of available 3D assets; even with large datasets like Objaverse-XL [deitke2023objaverse-xl], they still suffer from inconsistent quality and missing textures; (2) current systems struggle with fine-grained style control, unable to maintain consistent artistic styles (e.g., anime, cyberpunk) across all objects in a scene; and (3) most methods focus exclusively on indoor environments, lacking the flexibility to generate open-world scenarios like fantasy landscapes or futuristic cities.
@@
Recent progress in vision-language models (VLMs) [llava, molmo] and 3D generative models [poole2022dreamfusion] presents an unprecedented opportunity to address these limitations. VLMs now possess sophisticated spatial reasoning capabilities [chen2024spatialvlm] that can parse complex scene descriptions and generate physically plausible arrangements of objects. Simultaneously, state-of-the-art 3D generative models like Hunyuan3D [hunyuan3d2025hunyuan3d2.1] can create high-quality, stylistically consistent assets from text or image prompts. However, these advances have not yet been effectively integrated into a unified 3D scene generation framework.
@@
Facilitated by 3D asset generative models and the inference capabilities of VLMs, Holodeck 2.0 can produce high-quality 3D scenes with internal stylistic coherence for both indoor settings (e.g., music studios, playrooms) and open domains (e.g., airship docks, courtyards), as shown in Figure 1. Meanwhile, Holodeck 2.0 can generate the same type of 3D scenes in different styles, such as realistic, cartoon, and steampunk. The objects in Holodeck 2.0 scenes exhibit high quality and closely match their descriptions, even when those descriptions are stylized, imaginative, and richly detailed, for example, “A cartoon-style snail-shaped snack booth offers glowing drinks” as shown in Figure 1. In terms of spatial reasoning, Holodeck 2.0 fully leverages the spatial understanding capabilities of VLMs, iteratively generating and applying spatial constraints to produce layouts that match both textual and visual descriptions.
We demonstrate the effectiveness of Holodeck 2.0 through a comprehensive evaluation that encompasses large-scale human assessments, CLIP-based perceptual score, physical plausibility measures, and spatial-semantic alignment scores. Across 70 test cases spanning both indoor and open-domain scenes, Holodeck 2.0 consistently outperforms the baselines, with its advantage being particularly pronounced in open-domain scenes.
@@
In addition, we demonstrate Holodeck 2.0’s practical applicability through direct integration with commercial game engines. We showcase how a complex museum scene, complete with artifacts such as the Temple of Dendur, Venus de Milo, and The Thinker, can be generated from detailed textual descriptions and seamlessly imported into Unreal Engine with full interactivity. This shows how Holodeck 2.0 can accelerate 3D content creation workflows for game development and other applications.
In summary, our paper contributes as follows:
@@
    We present a unified framework that uniquely combines VLM-guided scene decomposition with generative 3D asset creation, utilizing VLMs across visual reference generation, object parsing, spatial constraint generation, and iterative layout refinement.
@@
    We demonstrate state-of-the-art performance on both indoor and open-domain scene generation with consistent artistic styles (realistic, cartoon, anime, cyberpunk, etc.) through extensive human and automated evaluations.
@@
    We validate downstream applications through integrating with commercial game engines and demonstrate significant quality improvements over baselines across perceptual, physical, and spatial-semantic metrics.
@@
Figure 2: Overview of Holodeck 2.0. Given a text input, Holodeck 2.0 generates 3D scenes via three modules: (1) Scene Analysis Module takes text as input and outputs object properties in JSON format, along with a reference image, individual object images, and a background image; (2) Object Generation Module takes individual object images as input and outputs textured 3D assets; (3) Scene Generation Module takes text, reference image, and object properties as input, and outputs the final 3D scene, with an optional editing module for human feedback.
2 Related Work
@@
Language-Guided 3D Scene Generation. Early work in this area is primarily symbolic and rule‑based, requiring extensive engineering effort and exhibiting very limited generalization ability, for example, Wordseye [coyne2001wordseye] and other systems [chang2014learningspatial, ma2018language3d]. With the emergence of LLMs and VLMs, open‑vocabulary 3D scene generation has become possible, greatly enhancing the flexibility of scene creation [yang2024Holodeck, hu2024scenecraft, zhang2025scenelanguage, lin2025partcrafter, ccelen2024I-design, fu2024anyhome, ocal2024sceneteller, deng2025global-local-tree-vlm, sun2025hierarchically, sun20253d-gpt]. Among them, Holodeck [yang2024Holodeck] employs LLMs to get the object descriptions that should appear in a scene based on the input text, and to generate spatial constraints among those objects. SceneCraft [hu2024scenecraft] also leverages LLMs and VLMs in a similar way, but it generates Blender code and iteratively optimizes it to produce 3D scenes. However, both of them rely on CLIP [radford2021clip] to retrieve 3D objects from asset libraries such as Objaverse [deitke2023objaverse] and TurboSquid [c:TurboSquid], and the variety and quality of these assets somewhat limit their capability to generate open‑world 3D scenes. Currently, larger or better‑curated 3D asset libraries have emerged [deitke2023objaverse-xl, lin2025objaverse++]. However, they still suffer from overall low asset quality, missing textures, and limited variety and styles of objects. For 3D scene spatial layout, LayoutGPT [feng2023layoutgpt] is a method that directly generates 3D object layout relationships via LLM prompting, but often neglects physical plausibility. LayoutVLM [sun2025layoutvlm] is another method that uses VLMs to generate 3D layouts by formulating objective functions that can be differentiably optimized. For 3D scene editing, recent works [c:BlenderGPT, huang2024blenderalchemy, gu2025blendergym] have applied LLMs or VLMs to editing, though they still face challenges in accurately managing object placement and layout.
@@
Text-to-3D Diffusion Models. The emergence of diffusion models [sohl2015deepthermo, ho2020ddpm, peebles2023dit, rombach2022stablediffu] marks a significant advancement in the field of generative modeling. Current methods leverage diffusion models to directly generate 3D objects or scenes represented by Neural Radiance Fields (NeRFs) or Gaussian splats from text or images [poole2022dreamfusion, lin2023magic3d, li2024dreamscene, po2024compositional]. However, such methods generate scenes that lack discrete, manipulable objects, making them difficult to edit directly and unsuitable for applications like game design or embodied AI training. When focusing solely on generating individual 3D assets, diffusion models offer significant advantages. Recent methods have achieved breakthroughs in generating high‑quality 3D assets [xiang2025TRELLIS, zhang20233dshape2vecset, zhang2024clay, zhao2023michelangelo, li2024craftsman3d, ren2024xcube, wu2024direct3d], with Hunyuan3D [hunyuan3d2025hunyuan3d2.1, lai2025hunyuan3d2.5] being a typical example.
3 Holodeck 2.0
Holodeck 2.0 is a vision-language-guided system, facilitated by VLMs and 3D generative models, that can produce and edit diverse, customizable, and stylistically varied 3D scenes.
As shown in Figure 2, Holodeck 2.0 systematically generates 3D scenes from text utilizing the following specialized modules: (1) Scene Analysis Module parses the object properties and background information that should appear in the scene from input text; (2) Object Generation Module leverages 3D asset generative models to produce 3D assets from quality‑controlled object images; (3) Scene Generation Module iteratively generates semantically and physically coherent object layouts based on spatial constraints; (4) Scene Editing Module automatically applies personalized adjustments to the scene based on human feedback.
In the following, we describe how each module of Holodeck 2.0 operates, specifically illustrating how it coordinates VLMs and 3D asset generative models to accomplish scene generation and editing. Detailed information about Holodeck 2.0 is also available in the supplement.
@@
3.1 Scene Analysis
The Scene Analysis Module, shown in the first panel of Figure 2, is designed to parse object properties that should appear in the scene from input text.
Text to Scene Image. Given a text input, the Scene Analysis Module uses GPT-Image-1 to generate a corresponding reference image. This image generation process is vital for the subsequent creation of uniform‑style assets. Moreover, visual inputs can substantially enhance the spatial awareness of large reasoning models. In Figure 3, we present several examples of scene reference images generated from text.
Scene Image to Object Images. After obtaining the scene image, the Scene Analysis Module employs GPT-o3 [GPTo3] to infer object properties (name, position, rotation, size, visual description, etc.) from the text input and its corresponding scene reference image. Then we prompt GPT-Image-1 to generate a clear, transparent frontal image of each object appearing in the scene image. This avoids the issues common to current image segmentation algorithms [kirillov2023segmentanything, ren2024groundedSAM], such as inaccurate segmentation, blurry edges, and difficulty handling occluded objects. After generating individual object images, we use GPT-o3 to perform quality control, removing smaller-object images redundantly embedded in larger-object images (e.g., removing the separate faucet when it is already present in the bathtub image). In practice, this approach generates accurate, high‑quality object images from the reference image as shown in Figure 3. Subsequently, we generate the scene’s background image based on the reference image, focusing primarily on the ground texture.
3.2 Object Generation
The Object Generation Module, shown in the second panel of Figure 2, utilizes 3D asset generative models to generate 3D assets from quality-controlled object images. For 3D asset generation, image-to-3D is faster and offers better style control than direct text-to-3D. After obtaining the object images from the Scene Analysis Module, we use a locally deployed Hunyuan3D 2.1 [hunyuan3d2025hunyuan3d2.1] to generate the corresponding 3D asset for each object. In practice, compared to searching for 3D objects in Objaverse, generative models can produce higher‑quality 3D assets that more faithfully match the descriptions (Figure 3), for example, “a plush, L-shaped sectional sofa adorned with colorful throw pillows” and “a cartoon-style donut-shaped bench”.
3.3 Scene Generation
Spatial Constraint Generation. The Scene Generation Module first employs GPT‑o3 to produce a series of spatial constraints between objects based on the text input, the scene reference image, and object properties. We define a total of ten spatial relations, which are divided into the following four categories: (1) Relative: in front of, behind, left of, right of, side of; (2) Distance: near, far; (3) Vertical: on, above and (4) Rotation: face to. Their mathematical definitions are provided in the supplement. As shown in Figure 4, reasoning models (e.g., GPT-o3) can generate realistic and spatially plausible constraints for objects.
Layout Generation. For layout generation, we employ a DFS solver to compute object placements based on object properties and spatial constraints between objects. We define mathematical formulations to enforce each spatial constraint (for example, “face to” means that the angle between the source object’s orientation and the line connecting the centers of the two objects is less than 10 degrees). Objects should satisfy these spatial constraints while avoiding physical collisions. To fully leverage GPT-o3’s spatial reasoning capabilities, we prioritize the initial object positions inferred by the model. It is worth noting that when the DFS solver encounters spatial constraints that are difficult to satisfy simultaneously, the Scene Generation Module re-prompts GPT-o3 using the DFS solver’s feedback to update the objects’ spatial constraints. The DFS solver and the reasoning model then form an iterative loop until a valid layout is found. This approach significantly enhances Holodeck 2.0’s layout capabilities. Figure 4 presents examples of position plots for the bounding boxes produced by the solver.
@@ 3.3 Scene Generation
3.4 Scene Editing
Scene Layout Level. If users want to personalize the generated scene layout after it has been created, they can provide a natural-language edit instruction. Then, it is incorporated into the original workflow to regenerate the spatial constraints between objects, and the DFS solver iteratively computes the updated layout.
3D Asset level. Utilizing 3D generative models, we support asset‑level editing. Because our scenes are built by assembling individual 3D assets, we can easily delete objects in response to user language feedback. Users can also replace objects, for example, “Change the torii gate to a pastel gradient color scheme and hang decorative bells on it.” The Scene Editing Module then uses that language feedback to regenerate a style‑consistent 3D asset for the specified object and substitute it into the scene. When adding new objects, the module regenerates suitable 3D assets and recalculates spatial constraints to maintain coherent scene layouts.